OpenStudy (explainitlikeimfive):

Solve by graphing. x^2 +2x - 8 = 0

OpenStudy (explainitlikeimfive):

–2, 4

OpenStudy (explainitlikeimfive):

2, –4

OpenStudy (explainitlikeimfive):

2, –4

OpenStudy (explainitlikeimfive):

–2, 4

OpenStudy (anonymous):

x^2+2x-8=0 (x+4)(x-2) x=-4, x=2 ans is b

OpenStudy (explainitlikeimfive):

Thanks